How to set up word problem: The measurement of the widthof a rectangle is the length decreased by 7 meters. The area is 60 square meters. Find the lenght and the width. I know the answer but not able to figure out how to set up the formula to get the right answer l= 12m & w= 5m. Thank You
Answer by mananth(16946)

length =x
width = x-7
Area = 60 m^2
Area = L*W
x(x-7)=60
x^2-7x=60
x^2-7x-60=0
x^2-12x+5x-60=0
x(x-12)+5(x-12)=0
(x-12)(x+5)=0
x=12 OR -5
ignore negative
so the length = 12 m
and the width is 7 less
so width is 5 m